OpenStudy (anonymous):

What is the general form of the equation of a circle with a center at (–5, 0) and a radius of 4?

OpenStudy (anonymous):

(x + 5)^2 + y^2 = 4 x^2 + (y + 5)^2 = 4 (x + 5)^2 + y^2 = 16 x^2 + (y + 5)^2 = 16

OpenStudy (anonymous):

\(\large (x-h)^2+(y-k)^2=r^2\) Center at (h,k) with radius, r.

OpenStudy (anonymous):

Thank you so much!

OpenStudy (aripotta):

what did you get for the answer?

OpenStudy (anonymous):

C , (x+5)^2+y^2=16

OpenStudy (aripotta):

yup. good job

OpenStudy (anonymous):

thank you!